OpenStudy (anonymous):

a flashlight has 6 batteries, 2 of which are defective, If 2 are selected at radom without replacement, find the proability that both are defective..........help

jimthompson5910 (jim_thompson5910):

P(Both Defective) = P(First defective AND second defective) P(Both Defective) = P(First defective) * P(Second defective) P(Both Defective) = (2/6) * (1/5) P(Both Defective) = 2/30 P(Both Defective) = 1/15

jimthompson5910 (jim_thompson5910):

Hopefully that all makes sense

OpenStudy (anonymous):

It sure does thank you for your help 5910

jimthompson5910 (jim_thompson5910):

yw